# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications

### v3

@@ 4.2.3. SwitchChannel
Broadcast. This primitive takes a source buffer as a local address on the local GPU and a destination buffer as a multimem address. The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes the multimem.st PTX instruction using the source register and the destination element’s multimem address. The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all the virtual memory addresses pointed by the multimem address.
4.3. MSCCL++ DSL API and Executor
MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level. It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back. MSCCL++ DSL is an extension of the MSCCLang (Cowan et al., 2023) that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time.
4.4. MSCCL++ Collective API and Kernels
MSCCL++ Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL API. We implement this API using kernels written in the MSCCL++ Primitive API. Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs. We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
@@ 4.4. MSCCL++ Collective API and Kernels
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficiently than other libraries by removing synchronizations.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collects and reduces $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) (Rabenseifner, 2004). Two-phase algorithms are more bandwidth-efficient and conduct less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple variants using PortChannel, MemoryChannel with LL or HB protocol, or SwitchChannel. Using MSCCL++, we can optimize 2PA in various ways that are not possible in existing libraries. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes each reduction step.
@@
4.5. Portability and Implementation Effort
MSCCL++ is designed to be portable. Code using MSCCL++ collectives, algorithms using the MSCCL++ DSL, and kernels using the MSCCL++ primitives are portable as long as the underlying MSCCL++ Primitive API supports that hardware. Our design makes it easy to add support for new hardware. As an example, MSCCL++ initially supported only NVIDIA GPUs; to add support for AMD MI300x GPUs, our implementation took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB. It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL. MSCCL++ could achieve this because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API. This significantly reduces the effort to maintain and develop MSCCL++.
The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features. For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast. Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $2.2\times$ in average for message sizes of 1KB - 1GB.
@@ 5. Evaluation
Environments. All presented numbers are collected from one of the environments listed in Table 1. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)), and inter-node IB links (one NIC per GPU). All NICs are connected to a single IB networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
Baselines. We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025a), RCCL 2.20.5 (AMD, 2025), [^5] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL. and MSCCL 2.23 (Microsoft, 2025a). MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms. Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
5.1. Collective Communication
To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 4.4 using the MSCCL++ DSL API. We also implement various AllGather algorithms in a similar manner, and we omit the implementation details. We present the best number among all implementations for each message size and environment. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025c)), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b). We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025f) and the CUDA/HIP Graph APIs for best performance.
For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^6] Defined as the message size divided by the latency.) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode (Pope et al., 2023)), while large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing a.k.a. prefill (Pope et al., 2023)).
Results. Figure 8 compares the AllReduce performance over A100-40G. MSCCL++ outperforms baselines for both small messages (up to 4.2x and 3.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.8x faster over both NCCL and MSCCL). Similarly, Figure 9 compares the AllGather performance over A100-40G. For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively. For AllGather, we observe a few cases where MSCCL slightly outperforms MSCCL++ by up to 8.0%, which is introduced by a performance bug of the DSL API, not by our algorithm implementation. For AllReduce, MSCCL++ is always the fastest.
@@ 5.1. Collective Communication
Gain Breakdown. The performance differences can come from both high-level differences in algorithms (i.e., design of data flow routes and pipelines) and low-level optimizations (i.e., how to transfer data, synchronize, and manage buffers). The former benefit is already delivered by MSCCL that enables the customization of algorithms over the NCCL communication stack. Therefore, the key benefits of MSCCL++ lie in the latter by exposing the primitive functionalities to users. Since MSCCL and MSCCL++ use the same high-level algorithm for all cases in Figure 8 and 9, the performance gap between the two comes purely from the benefits of MSCCL++.
@@
DSL vs Primitive. From our 9-month experience using the MSCCL++ DSL API, we find that it reduces development time from weeks to days, compared to using the Primitive API directly. The DSL helps automate initialization, check for mistakes, and provide performance analysis. We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL. However, as the DSL introduces higher level instructions than the Primitive API, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case. Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
Collective vs Primitive. For AI inference workloads, we implement the best algorithms in our collective kernels using the MSCCL++ Primitive API, so there is no performance difference.
Figure 10. Speedup for decodes with tensor parallelism = 8
5.2. LLM Inference
[…]

### v4

@@ 4.3. SwitchChannel
Broadcast. This primitive takes a source buffer as a local address on the local GPU and a destination buffer as a multimem address. The primitive goes through each element of the source buffer on the local GPU, reads the element into a register, and executes a multimem.st PTX instruction using the source register and the destination element’s multimem address. The multimem instruction sends the register value to the switch, which broadcasts and stores the value to all virtual memory addresses pointed by the multimem address.
4.4. Advantages of Primitive API
Expert developers can use the Primitive API to build custom workload-specific communication algorithms in a portable manner with minimal performance overhead (see Table 1). At the same time, developers implementing these primitives can add support for emerging hardware and interconnects. By separating primitives for data transfer and synchronization, MSCCL++ allows asynchronous communication and batching synchronization of multiple data transfers. By using channels for specific GPU interconnects, MSCCL++ allows users to specialize kernels for particular interconnects. Furthermore, MSCCL++ avoids extra copies in kernels, leading to a kernel with less code paths and zero register spills. Since the Primitive API is an in-kernel interface for communication, we can co-optimize both computation and the communication kernels, leading to better application performance. One example is overlapping reduction in AllReduce with communication (see Section 5.2).
Best Achievable MSCCL++
NVLink Throughput (GB/s) 397.5 397.5
NVLink Latency (ns) 822 829
InfiniBand Throughput (GB/s) 48.94 48.94
InfiniBand Latency ( µs) 3.76 4.89
Table 1. MSCCL++ Primitive API’s peer-to-peer performance compared to the best achievable on our H100 environment (see Table 2). The best achievable performance is measured by nvbandwidth (NVIDIA, 2025c) and RDMA perftest (RDMA, 2025). MSCCL++’s NVLink throughput and latency are measured with PortChannel and MemoryChannel, respectively. MSCCL++’s InfiniBand throughput and latency are measured with PortChannel.
5. MSCCL++DSL
1class PortChannel:
2 def __init__(self, dst_rank: int,
3 src_rank: int): ...
4 def put(self, dst_chunk: Buffer,
5 src_chunk: Buffer, tb: int): ...
6 def signal(self, tb: int): ...
7 def wait(self, tb: int): ...
8 def flush(self, tb: int): ...
9
10class MemoryChannel:
11 def __init__(self, dst_rank: int,
12 src_rank: int): ...
13 def put(self, dst_chunk: Buffer,
14 src_chunk: Buffer, tb: int,
15 tb_group: ThreadBlockGroup): ...
16 def put_packets(self, dst_chunk: Buffer,
17 src_chunk: Buffer, tb: int,
18 tb_group: ThreadBlockGroup): ...
19 def reduce(self, dst_chunk: Buffer,
20 src_chunk: Buffer, reduce_op: ReduceOp,
21 tb: int, tb_group: ThreadBlockGroup): ...
22 def signal(self, tb: int): ...
23 def wait(self, tb: int): ...
24
25class SwitchChannel:
26 def __init__(self, ranks: List[int]): ...
27 def broadcast(self,
28 dst_multimem_chunk: MultimemBuffer,
29 src_chunk: Buffer, tb: int): ...
30 def reduce(self, dst_chunk: Buffer,
31 src_multimem_chunk: MultimemBuffer,
32 reduce_op: ReduceOp, tb: int): ...
Figure 5. Channel abstractions in MSCCL++ DSL (Python).
To enable specifying custom communication algorithms, MSCCL++ supports a Python-based Domain-Specific Language (DSL) and lowers the algorithm into a sequence of operations that can be executed by the DSL Executor on top of the primitive abstractions described above. The primitives exposed at the DSL retain the zero-copy, one-sided, and asynchronous properties of the underlying primitive abstractions, thereby enabling users to specify optimized algorithms that overlap compute and communication.
1def ringRS(portChannels, bufs, sz, N, tb=0):
2 scr = [Buffer(rank, sz) for rank in range(N)]
3 for rank in range(N):
4 putChan = portChannels[(rank+1)%N]
5 recvChan = portChannels[(rank+N-1)%N]
6 src = bufs[rank]
7 recv, dst = scr[rank], scr[(rank+1)%N]
8
9 def chunk(step):
10 csz = sz / N
11 offset = ((rank + N - step) % N) * csz
12 return offset, offset+csz/2, offset+csz
13
14 for step in range(N):
15 beg, mid, end = chunk(step)
16 # (a) Put 1st half of chunk
17 putChan.put(dst[beg:mid], src[beg:mid], tb)
18 putChan.signal(tb)
19 if step != 0:
20 pbeg, pmid, pend = chunk(step-1)
21 # Reduce prev 2nd half, overlaps (a)
22 src[pmid:pend].reduce(recv[pmid:pend], tb)
23 recvChan.wait(tb) # Wait for 1st half
24 putChan.flush(tb) # Flush 1st half
25 # (b) Put 2nd half of chunk
26 putChan.put(dst[mid:end], src[mid:end], tb)
27 putChan.signal(tb)
28 # Reduce 1st half, overlaps (b)
29 src[beg:mid].reduce(recv[beg:mid], tb)
30 recvChan.wait(tb) # Wait for 2nd half
31 putChan.flush(tb) # Flush 2nd half
32 if step == N-1:
33 src[mid:end].reduce(recv[mid:end], tb)
Figure 6. A simplified example of a Ring ReduceScatter implementation using MSCCL++ DSL.
5.1. DSL
MSCCL++DSL provides a global view of all thread blocks on all GPUs or ranks in a Python-native language. As shown in Figure 5, the DSL exposes PortChannel, MemoryChannel, and SwitchChannel abstractions in this global view. Consequently, unlike the Primitive API (Section 4), users specify the ranks involved in each channel and for each operation in the channel, specify the thread block on the channel’s source rank that must execute it. Some operations on the MemoryChannel also accept an optional ThreadBlockGroup that enables a group of thread blocks to collectively perform that operation. The Channel classes include a few fused primitives (omitted in Figure 5 for brevity) that can reduce the overhead of API calls. For example, MemoryChannel’s reduce_put reduces two local buffers and puts the result to a remote destination using a temporary register to avoid a memory round-trip for the intermediate data. In addition, the DSL provides a separate Buffer abstraction and the Channels are not tied to a pair of buffers, so users specify the chunks of the source and destination buffer in each data transfer operations. Since the DSL is Python-native, the chunks are specified as slices of Buffer. Buffers are associated with a rank and allow computation like reduce, which is performed on its rank. The DSL also provides convenience classes to create input and output buffers based on the communication pattern (e.g., AllReduce, AllGather, ReduceScatter), device-wide synchronization primitives, and multi-device synchronizations primitives; synchronization within a thread block is handled automatically by the DSL. Users can control chunk sizes, number of queue pairs (QPs), etc. via the DSL. For instance, if users want two QPs, they can use two parallel PortChannels. If users want to offload the communication to CPU, they can use PortChannel rather than MemoryChannel. Thus, the MSCCL++ DSL enables users to construct complex data movement and synchronization workflows without writing low-level CUDA code.
5.2. Example: Overlapped Ring ReduceScatter
Figure 6 demonstrates how the DSL can implement an efficient Ring ReduceScatter algorithm that overlaps computation and communication. In the classic Ring algorithm, $N$ GPUs are organized in a ring, the input buffer is divided into $N$ chunks, and each chunk circulates around the ring to be reduced. At each step, a GPU receives a chunk, reduces it locally, and then forwards it to the next GPU, with all steps occurring synchronously. To enable overlap, Figure 6 further splits each chunk into two halves. This allows the reduction of the first half (Line 29) to overlap with the communication of the second half (Line 26), and the reduction of the second half (Line 22) to overlap with the communication of the first half (Line 17). In Figure 6, only thread block 0 on each GPU does this computation and communication. The user can experiment with more parallelism or pipelining by using more thread blocks on each GPU and/or using more chunks per GPU. The MSCCL++ DSL supports replicating program instances to increase parallelism and improve performance while automatically handling mapping of the program to thread blocks and channels.
5.3. Lowering and Optimizations
The program written in MSCCL++ DSL is lowered to an execution plan, which consists of information from the program such as the collective, channel types, data transfer protocol, memory buffers to register for each thread block, semaphores to setup, and the sequence of operations (which may include loops) to run in each thread block of every rank. Buffer sizes and rank counts (N in Figure 6) are provided by the user at lowering time. During lowering, the program goes through data dependence analysis to add appropriate synchronizations, as well as optimization passes such as operation fusion. This enables the DSL to achieve performance comparable to hand-tuned CUDA implementations.
Data Dependence Analysis. MSCCL++ DSL automatically tracks data dependencies at the chunk level within each thread block by maintaining the last writer and active readers for each memory slot. When operations have data dependencies, the lowered program includes necessary synchronization points to ensure correct execution order. For example, in Figure 6, it introduces a thread block synchronization on thread block 0 before the reduce operation (Line 29). The DSL also detects redundant synchronizations and removes them. Specifically, if a lowered program will be calling multiple thread block synchronizations back-to-back, the redundancies will be removed, retaining only one of them.
Operation Fusion. MSCCL++ DSL also maintains a directed acyclic graph (DAG) to track data dependencies and usage patterns of operations in a thread block at the chunk level. When two or more operations meet the fusion criteria, such as contiguous chunk access, no intervening dependencies, and compatible resource requirements, the DSL merges them into a single operation function. This fusion strategy reduces memory traffic and avoids unnecessary synchronization, resulting in faster execution. For example, a DSL code using MemoryChannel:
src.reduce(data,tb); memChan.put(dst,src,tb);is captured into a fused operation:
memChan.reduce_put(dst,src,data,tb);where reduce_put reduces two local buffers and puts the result to a remote destination using a temporary register to avoid a memory round-trip for the intermediate data.
@@ 5.3. Lowering and Optimizations
5.4. Executor
Given an execution plan and the input and output buffers, the DSL Executor initializes different channels, setting up connections and semaphores, and registering memory. Finally, the executor runs an execution kernel that interprets the list of operations in the execution plan. The execution kernel is a single GPU kernel implementation that runs any given execution plan. The execution kernel inlines calls to primitive MSCCL++ operations such as put, signal, wait, and flush. The execution plan can also contain fused operations, such as ReduceSend explained in Section 5.3.
6. MSCCL++Collectives Library
MSCCL++Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL or RCCL API. We implement this API using kernels written in the MSCCL++ DSL. Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs. We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficiently than other libraries by relaxing synchronizations and concurrently transferring data to multiple devices before waiting on any of them.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collects and reduces $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) (Rabenseifner, 2004). Two-phase algorithms are more bandwidth-efficient and conduct less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple variants using PortChannel, MemoryChannel with LL or HB protocol, or SwitchChannel. Using MSCCL++, we can optimize 2PA in various ways that are not possible in existing libraries. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes at each reduction step.
3. Two-phase Ring (2PR). Similar to 2PA, 2PR has two phases: the first phase does ReduceScatter in a ring and the second phase does AllGather in the same ring. Unlike NCCL, we can use PortChannel (DMA-copy) even within a node and we overlap the reduction with the DMA-copy. Figure 6 shows a simplified kernel for ReduceScatter. We extend such pipelining to overlap reduction with communication across both phases. Our 2PR implementation using PortChannel shows the best throughput among all implementations for intra-node AllReduce with large message sizes.
@@ 7. Evaluation
Environments. Table 2 lists the environments used in the evaluation. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300x) with intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)) and inter-node IB links (one NIC per GPU). All NICs are connected to a single IB networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
Baselines. We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025d), RCCL 2.20.5 (AMD, 2025), and MSCCL 2.23 (Microsoft, 2025a).
@@
7.1. Collective Communication Results
To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 6 using the MSCCL++ DSL. We also implement various AllGather algorithms in a similar manner, and we omit the implementation details. We present the best number among all implementations for each message size and environment [^3] We did not do sophisticated autotuning beyond picking the best configuration for a given platform and input size ranges (akin to MSCCLang (Cowan et al., 2023)) by offline profiling. Our experiments measure the performance of the resulting system. Orthogonal techniques from TACCL (Shah et al., 2023) and TE-CCL (Liu et al., 2024) can be used for automation, which is an interesting future direction.. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025a)), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b). We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025h) and the CUDA/HIP Graph APIs for best performance.
For clarity of exposition, we separate message sizes into small (up to 1MB, presented as latency) and large (1MB and above, presented as algorithm bandwidth (AlgoBW)) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode (Pope et al., 2023)), while large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing a.k.a. prefill (Pope et al., 2023)).
Results. Figure 7 compares the AllReduce performance over A100-40G. MSCCL++ outperforms baselines for both small messages (up to 4.2x and 3.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.8x faster over both NCCL and MSCCL). Similarly, Figure 8 compares the AllGather performance over A100-40G. For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively. MSCCL++ is faster than other baselines across the board for AllReduce. For AllGather, MSCCL outperforms MSCCL++ by 8.0% in one scenario. This is due to the performance overhead of our DSL API that we are able to reduce if we implement directly on top of MSCCL++’s Primitive API.
Gain Breakdown. The performance gain of MSCCL++ comes from both algorithmic improvements, fine-grained overlapping of compute and communication, and the ability to exploit low-level optimization capabilities provided by the Primitive API.
@@ 7.1. Collective Communication Results
DSL vs Primitive. From our 9-month experience using the MSCCL++ DSL API, we find that it reduces development time from weeks to days, compared to using the Primitive API directly. The DSL helps automate initialization, check for mistakes, and provide performance analysis. We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL. As the DSL introduces a runtime interpreter executor, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case. Despite performance overheads, MSCCL++ DSL is useful for quick prototyping of collective communication with easy-to-understand algorithm description.